did i solve this proportion correctly?
(2x+4)/5=30/25
(25)(2x+4)=(5)(30)
50x+29=150
50x+29-29=150-29
50x=121
(50x)/50=(121)/50
x=2.42

Found 2 solutions by Alan3354, oscargut:
Answer by Alan3354(69443)

did i solve this proportion correctly?
(2x+4)/5=30/25
(25)(2x+4)=(5)(30)
-------------
25*(2x+4) is 50x + 100, not 29. You would have gotten it right, even tho you went the long way around, except for that.
-------------------------
50x+29=150
50x+29-29=150-29
50x=121
(50x)/50=(121)/50
x=2.42
------------------
No.
(2x+4)/5=30/25 = 6/5
2x+4 = 6
2x = 2
x = 1
I don't know why you call it a "proportion", it's an equation.



Answer by oscargut(2103)

(2x+4)/5=30/25
(25)(2x+4)=(5)(30) (good)
50x+29=150 ( 25*4 =100 ) 50x+100=150
50x = 150-100
50x=50
x=1